OpenStudy (anonymous):

a piece of metal has a mass of 500 g and a volume of 5 cm^3. What is its specific gravity?

OpenStudy (anonymous):

it 7500. all u had 2 do was. times the 500*5*3= 7500. i did that 2 day in school. i know what it is. i new it already. no one in my class new it. but me. so if u need my help im here k. my email is amber.franklin5490@kirtland.edu k. if u need any help im here k. just email me and i will get back at u as soon as i can k. so i am looking forword from hearing from u. well have 2 go. i hope u have 2 great day. and have a nice weekend. love amber franklin class of 2012

OpenStudy (anonymous):

500/5=100 gr/cm3

OpenStudy (anonymous):

density of the metal = mass/volume =0.5/(5*10^-6) =10^5 kg m^-3 Specific gravity =density of the metal/density of water at 4 degree celcius =(10^5)/(10^3) =100

OpenStudy (anonymous):

A piece of metal has a mass of 500 g and a volume of 5 cm3. What is its specific gravity? 0.01 5 100 500 It =100. :)